# Good things come in small packages: Should we adopt Lite-GPUs in AI infrastructure?

### v1

@@ Abstract.
To match the blooming demand of generative AI workloads, GPU designers have so far been trying to pack more and more compute and memory into single complex and expensive packages. However, there is growing uncertainty about the scalability of individual GPUs and thus AI clusters, as state-of-the-art GPUs are already displaying packaging, yield, and cooling limitations. We propose to rethink the design and scaling of AI clusters through efficiently-connected large clusters of Lite-GPUs, GPUs with single, small dies and a fraction of the capabilities of larger GPUs. We think recent advances in co-packaged optics can be key in overcoming the communication challenges of distributing AI workloads onto more Lite-GPUs. In this paper, we present the key benefits of Lite-GPUs on manufacturing cost, blast radius, yield, and power efficiency; and discuss systems opportunities and challenges around resource, workload, memory, and network management.
@@ 1. Introduction
Demand for AI is growing and expensive to support (Minaee et al., 2024). These challenges are expected to only get harder as the diversity, complexity, and scale of AI models are growing, making it crucial for AI service providers to build powerful and efficient AI infrastructure (ope, 2024). However, scaling AI infrastructure is encountering significant obstacles (Morales, 2024). We have already reached the limit on how big a compute die can get, leading GPU designers to focus on advanced packaging technologies to pack more transistors into the same package (Figure 1). However, scaling an individual GPU package is becoming less and less sustainable for manufacturing due to multiple reasons, including power (Tirumala and Wong, 2024), cooling (Heydari et al., 2022), yield (Gupta and Lathrop, 1972; Teets, 1996), packaging costs (Tang and Xie, 2022), and failure blast radius (Liu et al., 2023a). For instance, the latest generation of NVIDIA GPUs is facing deployment delays due to packaging and cooling issues (Team, 2025; Garrou, 2024).
@@
We observe that there is an exciting alternative approach to scaling AI clusters. What if we replace large, powerful GPU packages with highly-connected clusters of Lite-GPUs, that each have only a single, smaller compute die and fractional performance? Smaller GPUs present many promising hardware characteristics: they have much lower cost for fabrication and packaging, higher bandwidth to compute ratios, lower power density, and lighter cooling requirements. In addition, they can also unlock desirable systems opportunities such as improved fault tolerance and finer-grained resource allocation. To date, distributing AI workloads to large number of GPUs has been challenging due to the data flow demanding very high-bandwidth communication across GPUs (Xu et al., 2021). Nevertheless, driven by recent advances in co-packaged optics, in the next decade, we expect off-package communication bandwidth to improve by 1–2 orders of magnitude with much better reach (10s of meters), compared to copper-based communication (Minkenberg et al., 2021; Tan et al., 2023; Zhang and Jia, 2022). Co-packaged optics intergrates electronic and optical components within millimeters, compared to current pluggable optics, cutting signalling distance and yielding better power efficiency. While there are open questions and active research on utilizing co-packaged optics, we think that it has the potential to disrupt the trade-off space around designing AI infrastructure, enabling Lite-GPUs that are equipped with high-bandwidth and energy-efficient optical interconnects, to communicate with many far-off Lite-GPUs at petabit per second bandwidths (Minkenberg et al., 2021; Tan et al., 2023).
@@ 2. The Lite-GPU
Through Lite-GPUs, we propose an alternative way of scaling AI clusters: with smaller but more GPUs connected through a performant and scalable network, realized through co-packaged optics. A Lite-GPU features a single compute-die GPU package where the die area is much smaller than that of state-of-the-art, leading to several hardware benefits. Figure 2 gives an example of Lite-GPU system where each NVIDIA H100 GPU is replaced with four Lite-GPUs. In this paper, we use this example for discussion and analysis that project future benefits of Lite-GPUs in AI clusters.
@@
Overall, we expect the operational cost of Lite-GPUs to be substantially lower due to better hardware yield and lower packaging cost. While the cost of networking should increase, we expect the net gains to be positive as the networking cost is only a small fraction of the GPU cost today. Additionally, there are many active efforts to scale networking cost sublinearly with network size using circuit switching (Ballani et al., 2020; Jouppi et al., 2023), which would allow for even larger Lite-GPU clusters.
@@ 3. Systems opportunities
As highlighted in the previous section, Lite-GPUs offer many hardware benefits and they can unlock the path towards efficient and scalable AI clusters. Nevertheless, some key research questions should be addressed so that we can realize the Lite-GPU disruption.
@@
Some of the research questions around using Lite-GPUs are not new or unique, but potentially amplified. For example, Lite-GPUs would result in more distributed systems in the datacenter, e.g., small models previously served by a single GPU are now distributed over multiple Lite-GPUs. For larger models that already need multiple GPUs, the number of devices would be multiplied. These can potentially amplify issues such as synchronization and straggling GPUs. Nevertheless, distributed ML training and inference is already in practice, and there is a lot of focus in having efficient and robust distributed AI platforms (Shoeybi et al., 2020; Kwon et al., 2023; Aminabadi et al., 2022).
@@
With Lite-GPUs, we can allocate and access smaller units of compute and memory, leading to greater flexibility in managing an AI cluster. For example, consider power management. A GPU’s compute clock frequency can be dynamically tuned to lower power consumption (Qiu et al., 2024). However, the granularity of down-clocking is on all Streaming Multiprocessors (SMs). SMs are processors designed for efficient parallel processing and each GPU consists of multiple SMs, similar to cores in a CPU. Down-clocking all SMs of a large GPU can lead to wasted resources or suboptimal performance. In a Lite-GPU cluster, we can control down-clocking at finer granularity to achieve better power efficiency, akin to down-clocking only a portion of SMs in a larger GPU.
Another example is around GPU configuration. Note that today, AI clusters with heterogeneous GPUs are already used to serve requests as power-efficiently as possible, e.g., by deploying different phases of transformer inference on different GPU hardware (Patel et al., 2024). We can customize and deploy Lite-GPUs for different profiles of AI workloads, similar to Splitwise, but at much finer scale, e.g., racks of custom Lite-GPUs as opposed to clusters of custom racks. Also, Lite-GPUs can allow for both easier over-clocking and higher bandwidth-to-compute ratios, potentially achieving higher performance efficiency at cluster-scale (Piga et al., 2024; Stojkovic et al., 2024b).
Third, these smaller GPU units may assist future AI as a service offerings. The ability to allocate small customizable Lite-GPU clusters per customer, separated physically, providing isolation and security, can be quite powerful.
@@
Careful workload parallelisation, deployment, and scheduling is a must in order to obtain the benefits of Lite-GPUs and to mask their overhead.
Most importantly, with Lite-GPUs, we move previously in-silicon traffic to optical network, potentially inducing additional latency and network load. Nevertheless, with AI workloads, there are several techniques we can use.
@@
Reducing the size of the GPU naturally reduces the blast radius should a GPU fail due to excessive temperatures, dust or debris, or transistor faults; leading to higher available FLOPS, memory capacity, and memory bandwidth at any time. However, the number of GPUs in the cluster is increased, potentially leading to different failure profiles. To maximize the benefit from smaller blast radii, building a robust and efficient software stack is crucial. Note that today’s large-scale inference pipelines already impose larger blast radii than the hardware-imposed blast radii: if one GPU out of group of GPUs serving a model instance fails, the entire instance is taken offline (Jouppi et al., 2023). Active work on resolving this issue can also help with Lite-GPU clusters (Strati et al., 2024; Miao et al., 2024).
@@
With regards to building an efficient, high-bandwidth Lite-GPU network, we have several options. First, as the traffic across Lite-GPUs that replace one large GPU is predictable, we can build a direct-connect topology within that group of Lite-GPUs and leave the remaining network as is. This is an approximation to the original network, though it eliminates the benefits of the smaller blast radius of Lite-GPUs. Alternatively, we can consider a (flat or hierarchical) switched network for the entire Lite-GPU cluster. Using circuit switching, in part or cluster-wide, may be crucial to achieve such a network at low cost. Circuit switching presents the following benefits over packet switching: (i) more than 50% better energy efficiency, (ii) lower latency, and (iii) more ports at high bandwidth, which allows for larger and flatter networks (Ballani et al., 2020).
@@
With Lite-GPUs the number of devices per area is increased, however, the energy per unit area is decreased. There is active research to handle data-center management at scale using various automation techniques which can be applicable to Lite-GPU clusters (Hong et al., 2024). Additionally, though the number of devices per rack may increase, the overall cooling requirements of the rack can be lighter due to the more efficient cooling of Lite-GPUs combined with co-packaged optics. This can eliminate the need for liquid cooling racks in the data-center, which comprise a significant portion of racks, and thus space, in an NVIDIA B200 cluster (nvl, [n. d.]).
@@ 4. Case Study: LLM Inference
(a) Prompt prefill. All configurations perform similarly. As the model sizes grow, the “Lite” cluster underperforms due to increased collectives causing network bottleneck. Increasing the network bandwidth compensates the increased network demand, overclocking improves performance further as prefill workloads are compute-bound.
(b) Decode. As model sizes and thus the number of required GPUs grow, the “Lite” cluster underperforms due to increased memory access intensities. The degredation is worse with GPT-3 due to it having more KV-heads resulting in propotionally longer memory-bound stages. As Lite-GPUs utilize their shoreline for more memory bandwidth, performance improves and exceeds the current H100 cluster.
@@
In this section, we present a case study of Lite-GPUs in the context of a trending AI workload –- LLM inference (Vaswani et al., 2023). LLM inference involves two distinct phases. The prompt prefill phase processes input tokens to compute reusable intermediate states, i.e., the Key-Value (KV) cache, and generates the first new token. The prefill phase is usually highly parallelizable and efficient in utilizing the compute resources. The decode phase generates output tokens one at a time, with each new token building on the entire KV cache and appending to it. This phase is often memory-bound and less efficient in compute utilization. In the evalulation, we assume that different phases can execute on different Lite-GPU clusters (Patel et al., 2024; Zhong et al., 2024) to demonstrate the hardware benefits that can be achieved with Lite-GPUs. With our case study on serving latest generative AI workloads, we aim to highlight future advantages of Lite-GPUs that are modified from today’s leading GPUs.
@@
We use roofline modeling (Williams et al., 2009) to capture important hardware and software characteristics and model a Lite-GPU cluster running LLM inference. We model important metrics including FLOPS, memory access, and network traffic of collectives. The modeling measures compute stages individually, including projection, MLP, and fused FlashAttention (Shah et al., 2024). Compute, memory I/O, and network I/O can overlap within each stage and tensor parallelism is used to distribute execution within each cluster.
NVIDIA H100 is the baseline GPU for comparison (Corporation, 2022). An H100 cluster consists of one to eight H100 GPUs. Each H100 includes 132 SMs. The Lite-GPU is modeled based on H100 by reducing its capabilities to 1/4 of the original, denoted as “Lite” in Table 1. Accordingly, a Lite-H100 cluster can consist of one to 32 Lite-GPUs, to match the total maximum number of SMs of the H100 cluster. Recall that for Lite-H100, we expect the bandwidth-to-compute can increase to 2×2$\times$2 × of H100 and that it can deliver higher sustainable FLOPS due to improved cooling efficiency. To explore how these hardware improvements can impact performance, we further define customized Lite-GPUs for comparison, as denoted and summarized in Table 1, with changed parameters in blue and red color.
@@
We evaluate performance with three LLM models with different sizes and structures: Llama3-70B, GPT3-175B, and Llama3-405B (Meta, 2025; Brown et al., 2020). We define the search criteria based on Splitwise’s latency requirements, with TTFT (time-to-first-token) $\leq$ 1s and TBT (time-between-tokens) $\leq$ 50ms constraints (Patel et al., 2024). We set a constant prompt sequence length of 1500 tokens, the reported median size in a production workload for coding (Patel et al., 2024). The search sweeps all possible batch sizes and number of GPUs to find the configuration with the best performance efficiency, that is the highest throughput per SM (tokens/s/SM). Note that while we sweep up to the maximum number of GPUs per cluster as defined in Table 1, the search may return that running a model with less GPUs than maximum yields better throughput per SM.

### v2

# Good things come in small packages: Should we build AI clusters with Lite-GPUs?

@@ Abstract.
To match the blooming demand of generative AI workloads, GPU designers have so far been trying to pack more and more compute and memory into single complex and expensive packages. However, there is growing uncertainty about the scalability of individual GPUs and thus AI clusters, as state-of-the-art GPUs are already displaying packaging, yield, and cooling limitations. We propose to rethink the design and scaling of AI clusters through efficiently-connected large clusters of Lite-GPUs, GPUs with single, small dies and a fraction of the capabilities of larger GPUs. We think recent advances in co-packaged optics can enable distributing AI workloads onto many Lite-GPUs through high bandwidth and efficient communication. In this paper, we present the key benefits of Lite-GPUs on manufacturing cost, blast radius, yield, and power efficiency; and discuss systems opportunities and challenges around resource, workload, memory, and network management.
*equal contribution, {burcucanakci,junyili,xingbowu}@microsoft.com
journalyear: 2025copyright: ccconference: Workshop on Hot Topics in Operating Systems; May 14–16, 2025; Banff, AB, Canadabooktitle: Workshop on Hot Topics in Operating Systems (HOTOS ’25), May 14–16, 2025, Banff, AB, Canadadoi: 10.1145/3713082.3730390isbn: 979-8-4007-1475-7/2025/05
[…]
by-nc-nd
@@ 1. Introduction
Demand for AI is growing and expensive to support (Minaee et al., 2024). These challenges are expected to only get harder as the diversity, complexity, and scale of AI models are growing, making it crucial for AI service providers to build powerful and efficient AI infrastructure (ope, 2024). However, scaling AI infrastructure is encountering significant obstacles (Morales, 2024). We have already reached the limit on how big a compute die can get, leading GPU designers to focus on advanced packaging technologies to pack more transistors into the same package (Figure 1). Nevertheless, scaling an individual GPU package is becoming less and less sustainable for manufacturing due to multiple reasons, including power (Tirumala and Wong, 2024), cooling (Heydari et al., 2022), yield (Gupta and Lathrop, 1972; Teets, 1996), packaging costs (Tang and Xie, 2022), and failure blast radius (Liu et al., 2023a). For instance, the latest generation of NVIDIA GPUs is facing deployment delays due to packaging and cooling issues (Team, 2025; Garrou, 2024).
@@
We observe that there is an exciting alternative approach to scaling AI clusters. What if we replace large, powerful GPU packages with highly-connected clusters of Lite-GPUs, that each have only a single, smaller compute die and fractional performance? Smaller GPUs present many promising hardware characteristics: they have much lower cost for fabrication and packaging, higher bandwidth to compute ratios, lower power density, and lighter cooling requirements. In addition, they can also unlock desirable systems opportunities such as improved fault-tolerance and finer-grained, flexible resource allocation.
To date, distributing AI workloads to large number of GPUs has been challenging due to the data flow demanding very high-bandwidth communication across GPUs (Xu et al., 2021). Nevertheless, driven by recent advances in co-packaged optics, in the next decade, we expect off-package communication bandwidth to improve by 1–2 orders of magnitude with much better reach (10s of meters), compared to copper-based communication (Minkenberg et al., 2021; Tan et al., 2023; Zhang and Jia, 2022). Co-packaged optics integrates electronic and optical components within millimeters, compared to current pluggable optics, cutting signalling distance and yielding better power efficiency. While there are open questions and active research on utilizing co-packaged optics, we think that it has the potential to disrupt the trade-off space around designing AI infrastructure. Notably, in the most recent GPU Technology Conference (GTC), NVIDIA highlighted their advancements in co-packaged optics in their effort to massively scale AI infrastructure with much improved power efficiency (NVIDIA, [n. d.]). We think co-packaged optics can enable Lite-GPUs that are equipped with high-bandwidth and energy-efficient optical interconnects, to communicate with many far-off Lite-GPUs at petabit per second bandwidths (Minkenberg et al., 2021; Tan et al., 2023).
@@ 2. The Lite-GPU
Through Lite-GPUs, we propose an alternative way of scaling AI clusters: with smaller but more GPUs connected through a performant and scalable network, realized through co-packaged optics. A Lite-GPU features a single compute-die GPU package where the die area is much smaller than that of state-of-the-art, leading to several hardware benefits. Figure 2 gives an example of a Lite-GPU system where each NVIDIA H100 GPU is replaced with four Lite-GPUs. In this paper, we mainly use this example while discussing potential benefits of Lite-GPUs in AI clusters.
@@
Overall, we expect the cost of Lite-GPUs to be substantially lower due to better hardware yield and lower packaging costs. While the cost of networking should increase, we expect the net gains to be positive as the networking costs are only a small fraction compared to the GPU costs today. Additionally, there are many active efforts to scale networking costs sublinearly with network size using circuit switching (Ballani et al., 2020; Jouppi et al., 2023), which would allow for even larger Lite-GPU clusters.
@@ 3. Systems opportunities
While scaling out AI clusters using today’s GPUs with co-packaged optics is an option, as highlighted in the previous section, Lite-GPUs offer many hardware benefits over current GPUs. So in this paper, we focus on utilizing Lite-GPUs as they have potential to unlock the path towards more efficient and scalable AI clusters. Nevertheless, some key systems research questions should be addressed so that we can realize the Lite-GPU disruption.
@@
Some of the research questions around using Lite-GPUs are not new or unique, but potentially amplified. For example, Lite-GPUs would result in more distributed systems in the datacenter, e.g., small models previously served by a single GPU are now distributed over multiple Lite-GPUs. For larger models that already require multiple GPUs, the number of devices would be multiplied. These can potentially amplify issues such as synchronization and straggling GPUs.
AI clusters come at different scales for training and inference, with training clusters being orders-of-magnitude larger, e.g., 16,000 vs 8 GPUs for Llama 3.1 405B (Meta, [n. d.]; Face, [n. d.]). An inference cluster with Lite-GPUs at the reduction ratios we discussed in Section 2 is unlikely to have more components than a training cluster today and would be easier to realize without heavy innovation in distributing models. In general, building efficient distributed ML training and inference platforms is an active research area and such approaches would also benefit clusters with Lite-GPUs  (DeepSeek-AI et al., 2025; Shoeybi et al., 2020; Kwon et al., 2023; Aminabadi et al., 2022).
@@
With Lite-GPUs, we can allocate and access smaller units of compute and memory, leading to greater flexibility in managing an AI cluster.
For example, consider power management. A GPU’s compute clock frequency can be dynamically tuned to lower power consumption during idle periods or to match stragglers (Chung et al., 2024; Qiu et al., 2024). However, the granularity of down-clocking is on all Streaming Multiprocessors (SMs). SMs are processors designed for efficient parallel processing and each GPU consists of multiple SMs, similar to cores in a CPU. Down-clocking all SMs of a large GPU can lead to wasted resources or suboptimal performance. In a Lite-GPU cluster, we can control down-clocking at finer granularity to achieve better power efficiency, akin to down-clocking only a portion of SMs in a larger GPU.
Conversely, we can over-clock Lite-GPUs to achieve higher performance while serving peak workloads, since smaller die areas allow for easier cooling and higher clock frequencies. Alternatively, more Lite-GPUs can be utilized to satisfy the peak load, but with the additional power overhead due to increased networking. Detailed analysis on workload patterns and power modelling can help us determine the most power-efficient approach for serving typical and peak workloads with Lite-GPUs.
Another example of resource management is around GPU configuration. Note that today, AI clusters with heterogeneous GPUs are already used to serve requests as power-efficiently as possible, e.g., by deploying different phases of transformer inference on different GPU hardware (Patel et al., 2024). We can customize and deploy Lite-GPUs for different profiles of AI workloads, similar to Splitwise, but at much finer scale, e.g., racks of custom Lite-GPUs as opposed to clusters of custom racks. Also, Lite-GPUs can allow for both easier over-clocking and higher bandwidth-to-compute ratios, potentially achieving higher performance efficiency at cluster-scale (Piga et al., 2024; Stojkovic et al., 2024b).
Third, these smaller GPU units may assist future AI as a service offerings. The ability to allocate small customizable Lite-GPU clusters per customer, that are separated physically and provide isolation and security, can be quite powerful.
@@
Careful workload parallelisation, deployment, and scheduling is a must in order to obtain the benefits of Lite-GPUs and to mask their overheads.
Most importantly, with Lite-GPUs, we move previously in-silicon traffic to an optical network, potentially inducing additional latency and network load. There are workloads that would be challenging to distribute further using Lite-GPUs, such as workloads that introduce randomness and congestion to the network traffic. Nevertheless, with AI workloads, there are several techniques we can use.
@@
Reducing the size of the GPU naturally reduces the blast radius should a GPU fail due to excessive temperatures, dust or debris, or transistor faults; leading to higher available FLOPS, memory capacity, and memory bandwidth at any time.
To maximize the benefit from smaller blast radii, building a robust and efficient software stack is crucial. Note that today’s large-scale inference pipelines already impose larger blast radii than the hardware-imposed blast radii: if one GPU out of a group of GPUs serving a model instance fails, the entire instance is taken offline (Jouppi et al., 2023). Active work on resolving this issue can also help with Lite-GPU clusters (Strati et al., 2024; Miao et al., 2024). One approach to dealing with such rigid, software-imposed GPU configurations is to include hot spares, spare GPUs that can be activated to serve a model instance while recovering from a failure. Lite-GPUs can suit this approach particularly well as a cluster of Lite-GPUs are larger with each additional Lite-GPU being smaller and cheaper. This reduces the proportional overhead of including spare Lite-GPUs, though we still need a strategy for how to best utilize them during normal operation.
In general, Lite-GPUs can help improve fault-tolerance of AI infrastructure. Nevertheless, with Lite-GPUs, the number of GPUs in the cluster is increased and additional networking components may be necessary, potentially leading to different failure frequencies and profiles. A thorough analysis of failures and recovery schemes is necessary to ensure that the reduced blast radius of Lite-GPUs are utilized.
@@
Another potential approach is to use Lite-GPUs along with disaggregated memory (Maruf and Chowdhury, 2023). Disaggregated memory can be used to provide a larger memory pool for Lite-GPUs and to allow for more efficient memory sharing across Lite-GPUs, though it introduces additional complexity in memory management. Note that, combined with the finer-granularity of Lite-GPUs, an AI cluster with Lite-GPUs, co-packaged optics, and disaggregated memory can enable us to flexibly adjust the compute-to-memory and compute-to-network ratios per Lite-GPU in the cluster.
@@
With regards to building an efficient, high-bandwidth Lite-GPU network, we have several options. First, as the traffic across Lite-GPUs that replace one large GPU is predictable, we can build a direct-connect topology within that group of Lite-GPUs and leave the remaining network as is. This is an approximation to the original network, though it eliminates the benefits of the smaller blast radius of Lite-GPUs. Alternatively, we can consider a (flat or hierarchical) switched network for the entire Lite-GPU cluster, yielding flexibility and improved fault-tolerance. Using circuit switching, in part or cluster-wide, may be crucial to achieve such a network at low cost. Circuit switching presents the following benefits over packet switching: (i) more than 50% better energy efficiency, (ii) lower latency, and (iii) more ports at high bandwidth, which allows for larger and flatter networks (Ballani et al., 2020).
@@
With Lite-GPUs, the number of devices per area is increased, however, the energy per unit area is decreased. There is active research to handle data-center management at scale using various automation techniques which can be applicable to Lite-GPU clusters (Hong et al., 2024). Additionally, though the number of devices per rack may increase, the overall cooling requirements of the rack can be lighter due to the more efficient cooling of Lite-GPUs combined with co-packaged optics. This can eliminate the need for liquid cooling racks in the data-center, which comprise a significant portion of racks, and thus space, in an NVIDIA B200 cluster (nvl, [n. d.]).
@@ 4. Case Study: LLM Inference
(a) Prompt prefill. All configurations perform similarly. As the model sizes grow, the “Lite” cluster underperforms due to increased collectives causing network bottlenecks. Increasing the network bandwidth compensates the increased network demand, overclocking improves performance further as prefill workloads are compute-bound.
(b) Decode. As model sizes and thus the number of required GPUs grow, the “Lite” cluster underperforms due to increased memory access intensities. The degradation is worse with GPT-3 due to it having more KV-heads resulting in proportionally longer memory-bound stages. As Lite-GPUs utilize their available shoreline for more memory bandwidth, performance improves and exceeds the current H100 cluster.
@@
In this section, we present a case study of Lite-GPUs in the context of a trending AI workload –- LLM inference (Vaswani et al., 2023). LLM inference involves two distinct phases. The prompt prefill phase processes input tokens to compute reusable intermediate states, i.e., the Key-Value (KV) cache, and generates the first new token. The prefill phase is usually highly parallelizable and efficient in utilizing the compute resources. The decode phase generates output tokens one at a time, with each new token building on the entire KV cache and appending to it. This phase is often memory-bound and less efficient in compute utilization. In the evaluation, we assume that different phases can execute on different Lite-GPU clusters (Patel et al., 2024; Zhong et al., 2024) to demonstrate the hardware benefits achievable with Lite-GPUs. With our case study on serving latest generative AI workloads, we aim to highlight potential advantages of Lite-GPUs that are modified from today’s leading GPUs.
@@
We use roofline modeling (Williams et al., 2009) to capture important hardware and software characteristics and to model a Lite-GPU cluster running LLM inference. We model important metrics including FLOPS, memory accesses, and the network traffic of collectives. The modeling measures compute stages individually, including projection, MLP, and fused FlashAttention (Shah et al., 2024). Compute, memory I/O, and network I/O can overlap within each stage and tensor parallelism is used to distribute execution within each cluster.
NVIDIA H100 is the baseline GPU for comparison (Corporation, 2022). An H100 cluster consists of one to eight H100 GPUs. Each H100 includes 132 SMs. The Lite-GPU is modeled based on H100 by reducing its capabilities to 1/4 of the original, denoted as “Lite” in Table 1. Accordingly, a Lite-H100 cluster can consist of one to 32 Lite-GPUs, to match the total maximum number of SMs of the H100 cluster. Recall that for Lite-H100, we expect that bandwidth-to-compute can increase to 2×2$\times$2 × of H100 and that it can deliver higher sustainable FLOPS due to improved cooling efficiency. To explore how these hardware improvements can impact performance, we further define customized Lite-GPUs for comparison, as denoted and summarized in Table 1, with changed parameters highlighted in blue and red.
@@
We evaluate performance with three LLM models with different sizes and structures: Llama3-70B, GPT3-175B, and Llama3-405B (Meta, 2025; Brown et al., 2020). We define the search criteria based on Splitwise’s latency requirements, with TTFT (time-to-first-token) $\leq$ 1s and TBT (time-between-tokens) $\leq$ 50ms constraints (Patel et al., 2024). We set a constant prompt sequence length of 1500 tokens, the reported median size in a production workload for coding (Patel et al., 2024). The search sweeps all possible batch sizes and number of GPUs for each GPU type. Then, since different GPU types have different hardware capabilities, we normalize the throughput for each configuration using the number of SMs in that configuration. The resulting metric, throughput per SM (tokens/s/SM), represents the performance efficiency of that configuration. For each GPU type, we plot the configuration with the highest throughput per SM. Note that while we sweep up to the maximum number of GPUs per cluster as defined in Table 1, the search may return that running a model with less GPUs than the maximum yields better throughput per SM.